OpenStudy (anonymous):

Find the derivative of: y=squareroot((x^(2)+9)/(x+3) @Mathematics

OpenStudy (anonymous):

\[y=\sqrt{(x ^{2}+9)\div(x+3)}\]

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

I know it is getting me so confused

OpenStudy (anonymous):

it will be \[\frac{1}{2\sqrt{\frac{x^2+9}{x+3}}}\times \frac{d}{dx}\frac{x^2+3}{x+3}\]

OpenStudy (anonymous):

second part you get from the quotient rule

OpenStudy (anonymous):

and of course the first part is \[\frac{\sqrt{x+3}}{2\sqrt{x^2+9}}\]

OpenStudy (anonymous):

I think it the derivative is equal to zero

OpenStudy (anonymous):

i made a typo in the first line, second part should be \[\frac{d}{dx}\frac{x^2+9}{x+3}\] and that derivative is \[\frac{x^2+6x-9}{(x+3)^2}\] by the quotient rule

OpenStudy (anonymous):

you have to set this equal to zero??

OpenStudy (anonymous):

NOpe

OpenStudy (anonymous):

good

OpenStudy (anonymous):

factor x^2+6x+9 (x+3)^2

OpenStudy (anonymous):

So th etop and bottome cross each other out and it becomes 1

OpenStudy (anonymous):

no there is no \[x^2+6x+9\] in it. it is \[x^2+6x-9\]

OpenStudy (anonymous):

OH I wish it was +9

OpenStudy (anonymous):

Thanks for your help ;D